OpenStudy (anonymous):

a 12-ft-long ladder is leaning against a wall and makes an 80 degree angle with the ground.how high up the wall does the ladder reach, and how far is the base of the ladder from the base of the wall? round to the nearest inch.

OpenStudy (cwrw238):



OpenStudy (cwrw238):

sin 80 = h / 12 h = 12 sin 80 where h = height of ladder cos 80 = b / 12 b = 12 cos 80 where b = distance bottom of ladder from wall

OpenStudy (anonymous):

so it is 69.1

OpenStudy (cwrw238):

no - what is 69.1?

OpenStudy (anonymous):

the answer......

OpenStudy (cwrw238):

there are 2 answers h - height and b - the distance on ground None of them can be as high as 69.1

OpenStudy (anonymous):

ohhh im definately lost :/

OpenStudy (cwrw238):

h = 12 * sin 80 = 12 * 0.9848 = 11.8 ft = 11ft 10 ins you can get the values of sin and cos of an angle on ur calculator

OpenStudy (anonymous):

ok but i still dont uderstand how u set that problem up?